# From Ground to Sky: Architectures, Applications, and Challenges Shaping Low-Altitude Wireless Networks

### v2

@@ II-C Standardization for LAWN-NTN Integration
The 3GPP has recently taken initial steps to accommodate aerial and NTN in its standards. Release 17’s scope of NTN encompassed both satellites and airborne platforms, identifying aerial drones and HAPs as special use cases within the NTN framework. In parallel, the system architecture work in Release 17 defined new functions to support aerial drone identification, authentication, and authorization over cellular networks. Building on this foundation, Release 18 expanded NTN capabilities to improve coverage and performance and to enable closer integration with terrestrial networks. On the aerial drone side, Release 18 introduced explicit new radio (NR) enhancements to better accommodate drones as network clients. In addition, the new “Mobile IAB” work item in Release 18 defined procedures for integrated access and backhaul (IAB) nodes mounted on moving vehicles, such as drones. This allows a drone equipped with a 5G base station to dynamically provide coverage to users in a cell, effectively creating a low-altitude cell site that can relocate as needed. Despite the above progress, the current 3GPP specifications mainly focus on the data-plane of the LAWN, while leaving gaps across the other three functional planes. In particular, C2 signaling, cooperative sensing, and edge-cloud coordination remain either optional or entirely absent, preventing LAWN from operating as a fully integrated cyber-physical platform.
@@ III-A Enabling Technologies
III-A1 Integrated Sensing and Communications (ISAC)
@@
ISAC is a transformative technology that allows performing both wireless communication and environmental sensing simultaneously by using the same signaling, hardware platform, and radio resources[7]. Traditionally, sensing and communication have been treated as distinct functions. However, in a LAWN, where drones continuously exchange information while also perceiving the dynamic environments, ISAC offers a highly efficient solution. By eliminating the need for separate sensing and communication modules, ISAC reduces spectrum usage and hardware complexity of the aerial nodes in LAWNs. As an example, a drone transmitting signals to a ground node can also monitor the reflected echoes from objects to detect emerging obstacles or track other drones without switching to a dedicated radar or LiDAR system. The received echoes at the LAWN nodes are processed via radar processing, like matched filtering and beamforming, followed by detection and estimation. With multiple drones acting as sensors over a wide area, the LAWN yields a large aperture, improving angular resolution and imaging capabilities. ISAC also facilitates channel state information acquisition via sensing rather than classical pilot-based channel estimation, further enhancing the system efficiency. Nevertheless, integrating ISAC into LAWN nodes imposes challenges such as waveform, hardware, and antenna designs.
III-A2 Delay-Doppler Waveform
Delay-Doppler (DD) waveforms, including orthogonal time frequency space (OTFS) and its variants, were proposed in 2017 for reliable data transmission in highly dynamic wireless channels. Unlike conventional time-frequency (TF) domain modulation, OTFS modulates symbols in the two-dimensional DD domain[8]. Since the physical world merely experiences a slight change in a short time duration, it is expected that the DD parameters associated with the channel are almost unchanged during a short symbol interval. Relying on the DD channel representation, OTFS converts the time-varying channels into near time-invariant ones, providing significant resilience when the channels suffer from high Doppler shifts and rapid channel fluctuations. Equalization in this domain becomes simpler; even if the drone or the scatterers are moving, their effect is concentrated in a few DD bins over the packet duration. Furthermore, DD waveform leverages full path diversity (all reflected delays/Dopplers) to maximize throughput under motion. Importantly, the DD waveform inherently carries physical meaning: a symbol’s spreading in the time-frequency signal corresponds to probing a range of delays and Dopplers. Thus, a LAWN receiver can naturally extract range and velocity information of reflecting objects from the communication signal itself. However, the 2D transforms involved can be computationally intensive for large signal bandwidths or high-resolution delay-Doppler grids, which is challenging for maintaining low-latency links.
III-A3 Semantic Communications
Rather than accurate bit-level transmission, semantic communications move the focus to successful delivery of task-relevant meaning while discarding bit redundancy. Instead of reconstructing transmitted messages with minimal bit error, semantic communication systems seek to understand the context, intent, and importance of the transmitted information, transmitting only the elements that directly affect the receiver’s task or inference. In LAWNs, a semantic-aware node may extract high-level descriptions rather than raw sensor data and transmit only the essential semantic representations, as discussed in the interest-based image segment transmission scheme[9]. This greatly reduces bandwidth consumption and energy usage, which is especially valuable for power-constrained drones. For example, rather than streaming a full 4K video feed, a drone may send a concise semantic message such as “three pedestrians crossing in two seconds,” allowing other LAWN nodes to react without processing the raw visual data. Semantic awareness also allows smarter control decisions. The controller can prioritize flows based on semantic importance, dynamically allocating bandwidth where it yields most task gain. To deploy semantic communications in the LAWN, the resource-constrained drones must run relatively sophisticated AI to interpret raw data. Another challenge is trust and correctness since some details are intentionally lost. In critical applications, an incorrect or oversimplified semantic message could lead to wrong decisions.
III-A4 Artificial Intelligence and Large Language Models
@@
AI is indeed the backbone technology in enabling the autonomy and intelligence of the LAWN [10]. It is closely connected with the capabilities for perception, decision-making, and cooperative behavior across spatially distributed nodes. By deploying lightweight neural networks on drones or edge servers, a wide range of real-time tasks such as channel prediction, trajectory optimization, obstacle classification, and intent estimation can be supported. The drones can be trained to navigate through complex environments or select optimal relay positions through reinforcement learning algorithms. The emergence of large language models (LLMs) has offered a new dimension of cognitive functionality, enabling reasoning, instruction interpretation, and symbolic decision-making. When integrated into LAWNs, LLM can translate natural-language instructions into structured mission parameters, enabling intuitive command interfaces for non-expert users. In turn, the mission logs, sensor feeds, or visual data obtained by the LAWN can be summarized as natural-language reports for post-mission review. Furthermore, LLM may assist semantic communications within multi-agent coordination through shared symbolic understanding[10]. Ensuring real-time performance of AI algorithms is crucial. Techniques like model compression, quantization, hardware acceleration (using onboard GPUs or NPUs), and offloading computation to edge servers are all employed to make AI feasible for resource-constrained drones. Moreover, the LAWN should incorporate online learning or periodic retraining to ensure that models stay robust over time.
III-A5 Edge Intelligence
The LAWN architecture includes a hierarchical computing plane: drones handle time-critical, lightweight tasks; edge servers manage latency-sensitive, compute-heavy functions; and cloud nodes perform long-term analytics. This hierarchy enables dynamic task partitioning[6]. For example, a drone may offload an intensive task like high-resolution image recognition to an edge server, while keeping tasks like simple obstacle avoidance on-board. A key advancement is edge perception, wherein rich, multimodal sensor data from multiple drones is transmitted to edge nodes for high-level scene understanding. Edge servers run deep models to generate real-time situational awareness and distribute the perception results back to the swarm. This approach enables each drone to benefit from a global perception system while minimizing onboard processing and inter-drone communication. A significant challenge in achieving edge intelligence in LAWNs is the 3D mobility, which causes intermittent connectivity and demands seamless handoff of tasks across edge nodes.
III-A6 Fully-Actuated Control Systems
@@
III-B Potential Applications of LAWNs
Relying on the LAWN’s reconfigurable 3D architecture, it opens a wide range of potential applications[12]:
One of the most impactful applications is in precision agriculture, particularly in mountainous regions and terraced fields. Using the LAWN, swarms of drones equipped with multispectral sensors can collect high-resolution vegetation indices across large fields. This data can be transmitted in real-time via the data plane to edge servers or aggregated onboard for immediate decision-making. Based on this feedback, the spraying drones can apply fertilizers or pesticides with centimeter-level precision, significantly reducing resource waste and environmental impact. The sensing and control planes coordinate flight paths to optimize the coverage, making large-scale agricultural automation both scalable and energy-efficient.
In urban logistics and air mobility, the LAWN provides the necessary infrastructure for reliable aerial transportation. Autonomous delivery drones can navigate complex city environments using shared 3D maps and real-time air traffic data spread through the LAWN. For eVTOL passenger services, the control plane can ensure safe airspace sharing, with real-time geofencing and cooperative obstacle avoidance. In both cases, the network’s dynamic reconfigurability allows for quick adaptation to incidents such as road closures, emergency zones, or changing weather.
The LAWN also enhances public safety and disaster response capabilities. In emergency scenarios, such as wildfires or earthquakes, terrestrial communication infrastructure may be damaged or overloaded. The LAWN can be rapidly deployed to establish ad-hoc communication networks, extending coverage to first responders and victims. Simultaneously, thermal or multisensor drones can search for survivors and deliver critical supplies. By leveraging the computing plane, these operations can be coordinated in real time, even when cloud access is limited.
Environmental and infrastructure monitoring is another key potential application. The sensing plane of the LAWN can measure air quality or extract the heat signature of forest ecosystems. For infrastructure, drones can regularly inspect bridges, power lines, pipelines, and rail networks, detecting faults such as cracks or corrosion using high-resolution imaging or LiDAR. The control plane ensures safe and repeatable inspection trajectories, while the data plane allows fast reporting and maintenance scheduling.
Beyond these established domains, the LAWN opens doors to emerging applications. In wildlife conservation, drones can track animal movements across inaccessible terrain, helping monitor migration patterns. In maritime or riverine environments, LAWN nodes can coordinate with aquatic drones for hybrid air-water sensing missions. In the mining industry, drones can perform underground inspections and real-time hazard monitoring, dramatically improving operational safety and efficiency. In smart cities, the LAWN may integrate with IoT systems to support crowd monitoring, traffic surveillance, or pollution tracking, providing governments with more information for urban planning.
III-C Lessons Learned
@@
  • Real-Time Data Processing: Leveraging its integrated data and computing planes, the LAWN enables real-time collection and processing of high-resolution data, as seen in precision agriculture, thereby improving resource optimization and decision-making in dynamic environments.
@@
Figure 4: Case study of LAWN-assisted swarm coordination. A swarm of aerial drones aim to survey damage and locate survivors. The four planes of the LAWN collaborate and adapt to ensure continuous and autonomous swarm operation.
@@ IV-A System Setup
In the system setup, we consider that a ground LAWN node (e.g., mobile command center) is equipped with multi-band radio transceivers and edge servers. It maintains a ground link to surviving terrestrial networks, a sub-GHz link for C2 signal, and an LEO satellite link as a backup. The LAWN architecture manages drone operations through four integrated planes: control, data, sensing, and computing. Specifically, the control plane employs time-synchronization beacons from the ground node to to keep all drones’ control loops aligned. The data plane carries mission payloads, such as high-resolution images, live videos, and sensory readings, via D2G links back to the ground node. Then, updated flight plans or map data are transmitted to the drones via G2D links, as shown in Fig. 4. The DD waveform is adopted in the data plane to preserve reliability under kHz-level Doppler shifts[13].
The sensing plane fuses with sensors, e.g., camera, LiDAR, and IMU with ISAC echoes, continually mapping the terrain and detecting survivors or obstacles. In a cluttered disaster zone, drones use sensing data to avoid collisions. The sensing plane thus augments the control plane with environment data, critical when “rapid changes, such as LoS/NLoS transitions” occur. The computing plane allocates heavy data processing tasks across onboard processors as well as the edge servers at the ground LAWN node, which does not slow down the aerial drones. Assuming that the formation task is for the drones to form a desired pattern. The ground LAWN node initially commands a leader drone to an entry point. Then the LAWN control plane ensures the other drones to adjusts their flight to maintain the desired formation.
Critically, even if the LAWN ground node loses partial or complete external connectivity, the four planes adapt to ensure continuous and autonomous swarm operation. The D2D links are established between drones to form a cooperative mesh. D2D links share sensed data, e.g. local obstacle maps, among neighboring drones. For instance, if one drone’s view is blocked, it can get surrounding map data via D2D links. Each drone’s onboard computing resources allow lightweight AI agents , enabling the swarm to dynamically reconfigure formations and continue executing mission tasks independently, ensuring robust and resilient operation under all network conditions.

### v3

@@ II-C Standardization for LAWN-NTN Integration
The 3GPP has recently taken initial steps to accommodate aerial and NTN in its standards. Release 17’s scope of NTN encompassed both satellites and airborne platforms, identifying aerial drones and HAPs as special use cases within the NTN framework [18]. In parallel, the system architecture work in Release 17 defined new functions to support aerial drone identification, authentication, and authorization over cellular networks. Building on this foundation, Release 18 expanded NTN capabilities to improve coverage and performance and to enable closer integration with terrestrial networks. On the aerial drone side, Release 18 introduced explicit new radio (NR) enhancements to better accommodate drones as network clients [1]. In addition, the new “Mobile IAB” work item in Release 18 defined procedures for integrated access and backhaul (IAB) nodes mounted on moving vehicles, such as drones. This allows a drone equipped with a 5G base station to dynamically provide coverage to users in a cell, effectively creating a low-altitude cell site that can relocate as needed.
Despite this progress, current 3GPP specifications remain largely data-plane centric and do not yet provide native support for LAWN’s integrated control, sensing, and I&C planes. In particular, standard primitives for safety-critical C2 reliability and redundancy, cooperative perception message exchange and trust, and task/model continuity across edge nodes are still missing or only loosely defined, which limits LAWN from operating as a fully integrated cyber-physical platform.
@@ III-A Enabling Technologies
III-B Enabling Technologies
The realization of LAWNs relies on a diverse set of enabling technologies, spanning physical-layer waveform design, integrated sensing, intelligent control, and network-level intelligence, as illustrated in Fig. 3. While Sec. II outlines the functional-plane architecture, the following tools and protocols are essential for achieving mission-grade reliability, situational awareness, and agile autonomy under fast topology changes and severe resource constraints.
III-B1 Integrated Sensing and Communications (ISAC)
@@
ISAC is a transformative technology that enables wireless communication and environmental sensing using the same signaling, hardware platform, and radio resources [14]. In a LAWN, drones must continuously exchange information while simultaneously perceiving a rapidly changing environment. Therefore, ISAC offers an efficient path to reduce spectrum usage and hardware complexity. For example, a drone transmitting to a ground node can also process reflected echoes to detect obstacles or track other aerial agents, without switching to a dedicated radar or LiDAR module. Unlike classic RF sensing using a probing signal $x(t)$ in (1), ISAC also conveys data information in the transmitted signal $x(t)$. The received echoes $r(t)$ can be processed using classical radar operations such as matched filtering and beamforming, followed by detection/estimation. Moreover, with multiple drones acting as distributed sensors, a LAWN can form a virtual large aperture, improving angular resolution and imaging capability. ISAC can further assist channel state (CSI) acquisition through sensing-assisted approaches, reducing the reliance on frequent pilot transmission [25]. For example, twenty questions estimation frameworks [29] could be potentially leveraged for CSI estimation. Nevertheless, bringing ISAC into compact aerial platforms poses practical challenges in waveform/antenna co-design, hardware limitations, and interference management.
III-B2 Delay-Doppler Waveform
[…]
III-B3 Semantic Communications
Semantic communications shift the goal from accurate bit-level recovery to successful delivery of task-relevant meaning, intentionally discarding redundancy that does not affect the downstream task. Rather than reconstructing raw messages with minimal bit error, semantic systems interpret context, intent, and importance, and transmit only representations that materially influence the receiver’s inference or action. This design can be captured by a task-oriented objective
[…]
$\text{s.t.}\quad R\leq R_{\max},\;\;\Pr(\mathrm{unsafe})\leq\epsilon,$
where $s$ denotes the underlying task-relevant state, $\hat{s}$ is its reconstruction at the receiver, and $\mathcal{L}_{\mathrm{task}}(\cdot)$ measures task degradation such as missed detection cost or control-performance loss. In LAWNs, a semantic-aware node may extract high-level descriptions from raw sensor streams and transmit only essential semantic representations, as in interest-based image-segment transmission [4]. This can significantly reduce bandwidth consumption and energy usage, which is particularly valuable for battery-constrained drones. For example, instead of streaming a full 4K video feed, a drone may transmit a concise semantic report such as “three pedestrians crossing in two seconds,” enabling other nodes to react without processing raw imagery. Semantic awareness can also support smarter control and scheduling by prioritizing flows based on task importance. However, semantic compression typically requires onboard or edge intelligence for interpretation, and it raises trust and correctness concerns: in safety-critical missions, incorrect or over-compressed semantic messages may lead to poor decisions[13].
III-B4 Artificial Intelligence and Large Language Models
@@
AI is indeed the backbone technology in enabling the autonomy and intelligence of the LAWN [27]. It is closely connected with the capabilities for perception, decision-making, and cooperative behavior across spatially distributed nodes. By deploying lightweight neural networks on drones or edge servers, a wide range of real-time tasks such as channel prediction, trajectory optimization, obstacle classification, and intent estimation can be supported. The drones can be trained to navigate through complex environments or select optimal relay positions through reinforcement learning algorithms. The emergence of large language models (LLMs) has offered a new dimension of cognitive functionality, enabling reasoning, instruction interpretation, and symbolic decision-making. When integrated into LAWNs, LLM can translate natural-language instructions into structured mission parameters, enabling intuitive command interfaces for non-expert users[3]. In turn, the mission logs, sensor feeds, or visual data obtained by the LAWN can be summarized as natural-language reports for post-mission review. Furthermore, LLM may assist semantic communications within multi-agent coordination through shared symbolic understanding. Ensuring real-time performance of AI algorithms is crucial. Techniques like model compression, quantization, hardware acceleration (using onboard GPUs or NPUs), and offloading computation to edge servers are all employed to make AI feasible for resource-constrained drones. Moreover, the LAWN should incorporate online learning or periodic retraining to ensure that models stay robust over time.
III-B5 Edge Intelligence
Edge intelligence complements the I&C plane plane by enabling distributed inference and fusion close to the operational area. A typical hierarchy assigns time-critical, lightweight functions to drones, latency-sensitive and compute-heavy tasks to edge servers, and long-term analytics/model training to cloud nodes, enabling dynamic task partitioning [11]. For example, a drone may keep immediate obstacle avoidance on-board while offloading compute-intensive recognition or multi-view fusion to a nearby edge server. A simple latency model that highlights the feasibility of offloading is
$T_{\mathrm{off}}=\frac{b}{R}+\frac{c}{f_{\mathrm{e}}}+T_{\mathrm{return}},$ (3)
where $b$ is the transmitted feature/data size, $R$ is the achievable uplink rate, $c$ is the required compute cycles, $f_{\mathrm{e}}$ is the edge processing rate, and $T_{\mathrm{return}}$ accounts for result delivery and protocol overhead. This expression illustrates that offloading is attractive only when the link is sufficiently stable and the edge is sufficiently close both in geometry and network load. A key capability is edge perception, where multimodal sensing data from multiple drones is aggregated at edge nodes for high-level scene understanding. The edge then distributes compact perception outputs back to the swarm, allowing each drone to benefit from shared situational awareness while limiting onboard processing and inter-drone traffic. A central challenge is 3D mobility, which causes intermittent connectivity and demands seamless task/model continuity across edge nodes during movement and handover.
III-B6 Multi-Connectivity and Redundant Networking
Mission-grade LAWN operation requires negligible performance degradation under blockage, interference, and node failures. Multi-connectivity provides a practical path to reliability by maintaining parallel links, e.g., sub-6 GHz control anchor with mmWave/FR2 data boosting, or terrestrial plus satellite backup[17]. Combined with packet duplication, opportunistic relaying over the A2A mesh, and lightweight coding strategies for multicast/group-oriented traffic, multi-connectivity can reduce disruption during fast LoS/NLoS transitions and make-before-break handoffs. In addition, redundant networking supports safety-critical traffic isolation, ensuring that C2 and remote-ID signaling remain protected even when mission payload traffic surges.
III-B7 Fully-Actuated Control Systems
@@
III-B8 Wireless Power Transfer
Energy supply is one of the most critical constraints for LAWNs. Conventional drones require frequent landings for recharging, which limits mission continuity. Wireless power transfer (WPT) can enable in-flight or hovering refueling without physical contact [22]. For far-field WPT using directional microwave/mmWave beams, the harvested energy over a charging interval $T$ can be abstracted as
$E_{\mathrm{WPT}}=\eta\,P_{\mathrm{t}}\,G_{\mathrm{t}}G_{\mathrm{r}}\!\left(\frac{\lambda}{4\pi d}\right)^{2}T,$ (4)
where $\eta$ is the end-to-end power-transfer efficiency, $P_{\mathrm{t}}$ is the transmit power, $G_{\mathrm{t}}$ and $G_{\mathrm{r}}$ are transmit/receive beamforming gains, $\lambda$ is the wavelength, and $d$ is the separation distance. It can be observed that the harvested energy relies on geometry and beam alignment, implying that WPT must be co-designed with routing and formation control, e.g., by guiding drones into favorable charging locations and maintaining beams under mobility.
III-C Applications enabled by LAWNs
LAWNs unlock a broad range of applications across consumer, industrial, and public sectors by combining altitude-segmented air corridors with tightly coupled functional planes.
Urban logistics is among the most visible application domains for LAWNs, covering parcel delivery, food distribution, and medical supply transport. Vertiports, dynamic geofencing, and integrated traffic management within the control plane keep UAV logistics both efficient and compliant with safety regulations. This is particularly critical for last-mile delivery in dense urban environments and for rapid deployment of essential medical supplies during emergencies [8].
In rural and regional settings, LAWNs enable large-scale sensing and data collection for crop monitoring, spraying, irrigation planning, and environmental surveillance[23]. The sensing plane supports multimodal perception to produce high-resolution environmental maps. Data aggregation and processing can be offloaded to edge computing facilities, while long-horizon analytics such as yield prediction or climate impact modeling reside in the cloud. These systems combine centimeter-scale accuracy with edge analytics, providing timely warnings in remote regions.
In disaster and emergency scenarios, LAWNs offer a way to rapidly restore connectivity and generate situational awareness. When terrestrial infrastructure is damaged, drones can form airborne relay networks to re-establish communication links and provide real-time imagery for search and rescue. The control plane maintains robust C2 connections, while the sensing plane aggregates observations for localizing survivors and identifying critical hazards. Beyond acute emergencies, the same infrastructure can support routine public safety tasks, including traffic monitoring, crime prevention, and firefighting in dense urban areas [16].
Urban Air Mobility, powered by eVTOL aircraft and other emerging platforms, is one of the most transformative application families enabled by LAWNs. Over the past few years, eVTOL test flights and demonstrations have taken place across North America, Europe, and Asia [5]. Turning these demonstrations into scalable services will require low-altitude corridors with URLLC for safety-of-flight, cooperative sensing for collision avoidance, and seamless integration with UTM/U-space systems. LAWNs provide the necessary infrastructure for these requirements. The control and data planes jointly support fleet coordination, route scheduling, and integration of crewed and uncrewed traffic. The sensing plane supplies shared perception and traffic awareness, while the I&C plane handles trajectory optimization, demand-responsive routing, and predictive maintenance.
Infrastructure inspection is another high-impact use case in which LAWNs play a central role. drones can inspect bridges, pipelines, power lines, offshore wind farms, and other critical assets, often in environments with complex propagation conditions and limited ground access. Field reports indicate that major utilities in California have saved millions of dollars and improved worker safety by switching to drone-based inspections for power lines [6].
Beyond mission-critical domains, LAWNs also unlock new applications in entertainment and media [15]. Coordinated drone swarms have already demonstrated large-scale aerial light shows, surpassing the capabilities of traditional fireworks by offering programmable, environmentally friendly, and reusable visual displays. In addition to these established applications, LAWNs open the door to immersive AR/VR experiences in stadiums and city centers, where synchronized aerial platforms can project visual effects or serve as mobile anchors for mixed-reality overlays. Looking further ahead, LAWNs may support emerging services such as airborne data centers and persistent aerial sensing grids that integrate into smart-city infrastructures.
In summary, this flexibility of LAWNs ensures can they can simultaneously support mission-critical services like public safety while also fostering new commercial and consumer ecosystems in the low-altitude airspace.
III-D Lessons Learned
@@
  • Real-Time Data Processing: Leveraging its integrated data and I&C planes, the LAWN enables real-time collection and processing of high-resolution data, as seen in precision agriculture, thereby improving resource optimization and decision-making in dynamic environments.
@@
Figure 4: Case study of LAWN-assisted swarm coordination in post-disaster scenarios. A ground node disseminates flight/formation updates via G2A links and reliable C2 signaling, collects mission payloads via A2G links, and leverages A2A links among drones for local information sharing and cooperative coordination.
@@ IV-A System Setup
In the system setup, we consider that a ground LAWN node (e.g., mobile command center) is equipped with multi-band radio transceivers and edge servers. It maintains a ground link to surviving terrestrial networks, a sub-GHz link for C2 signal, and an LEO satellite link as a backup. The LAWN architecture manages drone operations through four integrated planes: control, data, sensing, and I&C. Specifically, the control plane employs time-synchronization beacons from the ground node to to keep all drones’ control loops aligned. The data plane carries mission payloads, such as high-resolution images, live videos, and sensory readings, via A2G links back to the ground node. Then, updated flight plans or map data are transmitted to the drones via G2A links, as shown in Fig. 4. The DD waveform is adopted in the data plane to preserve reliability under kHz-level Doppler shifts.
The sensing plane fuses with sensors, e.g., camera, LiDAR, and IMU with ISAC echoes, continually mapping the terrain and detecting survivors or obstacles. In a cluttered disaster zone, drones use sensing data to avoid collisions. The sensing plane thus augments the control plane with environment data, critical when “rapid changes, such as LoS/NLoS transitions” occur. The I&C plane allocates heavy data processing tasks across onboard processors as well as the edge servers at the ground LAWN node, which does not slow down the aerial drones. Assuming that the formation task is for the drones to form a desired pattern. The ground LAWN node initially commands a leader drone to an entry point. Then the LAWN control plane ensures the other drones to adjusts their flight to maintain the desired formation.
Critically, even if the LAWN ground node loses partial or complete external connectivity, the four planes adapt to ensure continuous and autonomous swarm operation. The A2A links are established between drones to form a cooperative mesh. A2A links share sensed data, e.g. local obstacle maps, among neighboring drones. For instance, if one drone’s view is blocked, it can get surrounding map data via A2A links. Each drone’s onboard computing resources allow lightweight AI agents , enabling the swarm to dynamically reconfigure formations and continue executing mission tasks independently, ensuring robust and resilient operation under all network conditions.